# What Sensors See, What People Feel: Exploring Subjective Collaboration Perception in Mixed Reality

### v1

@@ Abstract
Mixed Reality (MR) enables rich, embodied collaboration, yet it’s uncertain if sensor and system-logged behavioral signals capture how users experience that collaboration. This disconnect stems from a fundamental gap: behavioral signals are observable and continuous, while collaboration is interpreted subjectively, shaped by internal states like presence, cognitive availability, and social awareness. Our core insight is that sensor signals serve as observable manifestations of subjective experiences in MR collaboration, and they can be captured through sensor data such as shared gaze, speech, spatial movement, and other system-logged performance metrics. We propose the Sensor-to-Subjective (S2S) Mapping Framework, a conceptual model that links observable interaction patterns to users’ subjective perceptions of collaboration and internal cognitive states through sensor-based indicators and task performance metrics. To validate this model, we conducted a study with 48 participants across 12 MR groups engaged in a collaborative image-sorting task. Our findings show a correlation between sensed behavior and perceived collaboration, particularly through shared attention and proximity.
@@
Mixed Reality, Collaboration, PerceptionIntroduction
Collaboration within mixed reality (MR)111MR aligns with Milgram’s reality-virtuality continuum [33], closely related to Augmented Reality (AR), where virtual elements seamlessly integrate with the real world. environments is becoming increasingly prevalent across domains such as education, design, healthcare, and remote work [15]. This trend stems from MR’s unique capabilities to seamlessly integrate physical and virtual realms, enabling co-located and remote users to interact through shared digital content. However, as these systems evolve to become more intricate, immersive and integral to collaborative work, researchers have explored various aspects of MR collaboration, including the effects of gaze, gesture sharing, and virtual replicas on presence and cognitive load [4, 56], as well as domain-specific applications such as radiography training [64, 9].
Despite these advances, a fundamental challenge persists: How can we effectively understand the quality of collaboration that unfolds within these environments? Modern MR systems generate extensive behavioral data through system logs and embedded sensors, capturing where users look [44], how they move [29], and when they speak [48], yet these signals often provide only limited insights into users’ internal states [68]. The critical gap in current research is understanding whether such sensor data meaningfully reflect participants’ actual collaborative experiences or whether important dimensions of collaboration remain undetected by current measurement approaches.
This paper takes a human-centered approach to this question. Rather than treating collaboration as an externally defined construct, we focus on how participants themselves perceive their group interactions. In doing so, we recognize that collaboration is not merely a behavioral pattern but a subjective [38], socially mediated experience [42]. For MR systems to effectively support teamwork, we need methods that not only detect and interpret this experience but also bridge the gap between observable actions and users’ perceptions of their collaborative interactions.
To explore this, we study a collaborative MR task using a multi-perspective approach involving 48 participants grouped into 12 teams of four. First, we analyze sensor-derived indicators of group interaction, including gaze-based shared attention, conversation dynamics through audio, and spatial proximity, which can be captured through embedded sensing in MR headsets [68, 65]. Second, we examine task performance using system logs, measuring timing, interaction patterns, and decision changes throughout the collaboration. Finally, we assess subjective experience through post-exposure questionnaires that capture participants’ perceptions of group dynamics as well as their individual sense of presence and cognitive effort.
@@ 1.2 Sensor-Based Indicators of Group Behavior
Sensor-based methods across ubiquitous and immersive computing now characterize group behavior through multiple data streams (gaze, voice, motion, location) without relying on external observation or self-reporting. Remote collaboration systems, such as TeamSense [68] and CoCo [46] monitor nonverbal cues to track group cohesion, while wearable computing approaches assess synchrony and proximity patterns during collaborative tasks [54]. In immersive environments, headset-embedded sensors measure participation and spatial orientation, and researchers demonstrate how gaze and object manipulation reflect collaboration quality in VR tasks [65] and employ volumetric capture to assess gesture alignment in MR learning contexts [26, 22]. Despite these advances, most existing approaches emphasize either performance outcomes or structural interaction properties (such as coordination patterns and role distribution) without examining their relationship to participants’ subjective experiences. The critical question of whether sensor-based behavioral indicators meaningfully reflect users’ internal perceptions of collaboration remains largely unexplored, particularly in co-located MR tasks using commercial headsets.
@@ 1.3 Subjective Indicators of Group Behavior
Subjective self-report surveys remain central to understanding how people experience collaboration. Standardized tools, such as post-task surveys, presence questionnaires [62, 47], and cognitive load surveys [20], are widely used in the Human-Computer Interaction (HCI) literature to gauge how users evaluate, interpret, and internalize virtual experiences and their experience. Previous work has used these subjective measures to assess collaborative quality in various contexts, including team cohesion in distributed work [68], perceived social presence in remote collaboration [41], and mutual understanding during joint problem-solving [67]. Researchers have used self-report data in educational settings to evaluate perceived contribution equity and satisfaction with group outcomes [61]. Similarly, studies of collaborative design have incorporated subjective assessments of idea sharing and creative synergy [36].
Despite rich insights, these subjective measures are rarely systematically linked to sensor-based behavioral data. This disconnect raises questions about whether group behaviors in MR environments actually reflect users’ internal collaborative experience. If MR systems can detect where users are looking, who is speaking, and how close they are, can those signals be meaningfully interpreted as indicators of subjective collaborative quality? Do behavioral patterns track with subjective perception or challenge our assumptions about collaboration? Addressing these questions requires examining the relationship between sensor data and human experience in collaborative MR contexts.
@@ 2.1.1 Subjective Perception of Collaboration
Collaboration is not merely the accumulation of actions performed by individuals; it is shaped by how each user interprets, feels, and reflects on those interactions. Collaboration is what group members come to understand and feel about what they did together. In MR, where social cues, task progress, and attention are distributed across a blend of physical and virtual elements, this perception is especially complex. To understand collaboration in such settings, we must look not only at visible behavior but also at the subjective interpretations users form before, during, and after interaction. These interpretations are shaped by both individual-level experience and group-level awareness, which together constitute what we refer to in this paper as subjective perception of collaboration. As such, we treat the subjective perception of collaboration as a layered construct, one that blends individual experiences with group-level awareness. This section unpacks that construct and explains how we operationalize it in our study.
Individual-Level Experience. Each user in a collaborative MR task brings their own perspective on the interaction. This perspective is filtered through their sense of presence [50], cognitive availability [52], and personal contribution [12]. These factors are not passive background states but directly influence how users perceive and make sense of social interaction [41]. At its most foundational, collaboration begins with how individuals experience being part of a group in each moment. In immersive settings, this includes:
Presence: the degree to which a user feels immersed and situated (“being there”) in the virtual environment is the foundational phenomenon. Presence affects how responsive or attuned a user is to the behaviors of others. Prior research has shown that presence increases sensitivity to others’ gaze, gesture, and speech, and enhances the salience of social cues [49, 47]. In our context, a greater presence may facilitate awareness of others’ engagement, enabling more accurate reflections on group behavior.
Cognitive Load: often understood as the mental effort required to manage task complexity, modulates how much a user can track, interpret, and respond to others. A mentally taxed user may overlook subtle coordination signals, while a cognitively available user may notice shifts in attention, interruptions in conversation, or emerging group patterns [13]. In this way, we hypothesize that cognitive load shapes not just task performance [24], but the perceived quality of collaboration. Higher cognitive load may inhibit attention to group interactions or reduce perceived performance in the collaboration [55].
Contribution Awareness: how much a user feels other members contributed meaningfully to the group’s progress. This involves awareness of others’ actions and roles within the group. This acts as a bridge between personal effort and social perception. It reflects both internal judgment and observed input from peers. In immersive collaboration, where explicit roles may not be assigned, contribution awareness can become a key proxy for how individuals interpret participation and balance.
@@
Group-Level Reflection. While the MR system tracks each user individually, users themselves develop relational interpretations: Were we focused on the same thing? Did we support each other’s efforts? Did we function as a group? These are not directly observable but accessible through post-task subjective reflection [53]. We capture these interpretations through four key constructs:
Shared Attention Awareness reflects user’ recognition that others are attending to the same virtual elements. This perception is not only a sign of visual convergence, but it also reflects shared intention and awareness, which are foundational for coordination and meaning-making [34].
Conversational Support measures whether dialogue with others helped a user understand the task and contribute effectively. It captures the conversation and its function: Did it clarify? Did it invite participation? This reflects group cohesion and mutual scaffolding [18].
Proximity Impact probes how user felt physical closeness (as mediated by MR headsets) influenced collaboration. Prior work suggests that co-located interaction, being physically near others, improves responsiveness. [17].
Group Collaboration captures users’ holistic sense of how well the team worked together. This reflects an integration of observed behaviors, internal judgments, and implicit comparison to collaborative expectations [16].
Though collected individually, these measures represent users’ interpretations of the group. They are formed through the lens of each user’s own state, engagement, and task involvement. In other words, they are subjective readings of an inherently social phenomenon.
@@ 2.1.2 Objective Perception of Collaboration
While subjective reports offer insight into how collaboration is experienced, immersive systems provide a complementary channel: detailed logs of where users look, how they move, and when they speak. These objective and sensor-based indicators, derived from sensor data and system instrumentation, allow us to capture collaboration not as it is reported but as it unfolds. Although these are individual-level measurements, they are not isolated; they reflect what each user does in relation to others in the group. The same user performing the task alone would not produce the same sensor traces. These data are fundamentally shaped by interaction: who a user looks at, how close they are to others, and when they speak or remain silent. Each signal gains meaning only in the context of collective activity. The central challenge lies in interpreting these distributed signals in ways that meaningfully reflect group-level coordination. Our system logs three core types of interaction indicators, chosen for their theoretical grounding in studies of group behavior [65, 68] and practical availability from commodity MR headsets such as Meta Quest Pro [32], Hololens 2 [1], and Apple Vision Pro [2]:
Conversation and speech activity has been frequently employed as a key metric for examining behavior within small groups in the fields of psychology, anthropology, and traditional observational studies [28, 10]. Even recent work in the literature utilizing ubiquitous sensors to understand group behavior continues to emphasize the significance of this metric [27]. Studies have demonstrated that the proportion of conversation among humans varies across different media, such as virtual reality and desktop environments [65]. Given that MR immersive experiences offer a unique modality for human interaction with virtual objects while remaining grounded in the physical world, capturing conversation and speech metrics is valuable for gaining insights into group behavior. To this end, speech activity is logged via audio data collected from the microphones embedded in the MR headsets to capture conversation metrics. Rather than transcribing content, we focus on conversational metrics: speaking time distribution, number of turns, and participation variance across group members. These features have been shown to correlate with dominance and conversational fluency in other settings [37].
Shared attention, a cognitive process involving the ability of individuals to focus with a social partner on a shared object of interest for either intentional or social purposes, has been extensively studied in the literature [7, 34]. Various works in the literature have consistently highlighted the critical role of shared attention in facilitating cooperation and social bonding [63]. Given its pivotal role in collaborative behavior, the significance of shared attention in the context of computer-supported collaborative work has been a subject of considerable exploration [67]. Many studies have used gaze awareness for automatic shared attention detection. Building on this, we detect shared attention episodes by analyzing overlapping gaze vectors on the same virtual object using eye-tracking data from MR headset sensors. This operationalizes shared attention [34] in a continuously measured, time-resolved, and automatically logged way. While shared attention is inherently social, it is not explicitly communicated; users may not be aware of it. We aim to assess whether sensed shared attention reflects users’ perception of collaboration.
@@
Proximity plays a critical role in collaboration and group interactions. Prior work in organizational behavior has shown that closer physical proximity among team members enhances knowledge exchange, creativity, and teamwork quality [60]. In co-located collaborative scenarios, groups tend to perform better when members are physically near one another rather than dispersed [21, 59]. Additionally, studies have observed that individuals often move closer to those with whom they share stronger social bonds [14]. Building on these insights, we capture proximity as a scalar measure, specifically, the average pairwise distance between users, using positional data from each user’s headset in a shared coordinate frame.
@@
We deliberately abstract away from directionality in terms of proximity and conversation (e.g., who was closest to whom or who spoke to whom) and instead focus on group-level distributions. This is because our interest lies in capturing whether everyone engaged with everyone, not in identifying dominant or passive roles. Prior work has shown that proximity itself (regardless of direction) is a meaningful cue of social engagement in co-located settings [14]. Moreover, our subjective measures focus on perceived group collaboration, which is more likely influenced by general closeness than specific spatial arrangements. Prior work supports this approach, showing that evenly distributed proximity and turn-taking indicate group cohesion [68].
@@
In addition to sensor-based metrics, we log task-relevant interactions that reflect how users engage with the collaborative activity over time. These include interaction frequency, object manipulation patterns, and changes to previously completed actions. For example, we track how often each user interacts with shared content, how frequently prior actions are revisited or overridden, and the number of unique configurations a group explores before reaching a consensus. These metrics offer a view into participation distribution, coordination dynamics, and decision negotiation, which are key elements of collaborative processes. We also record overall task completion time, which serves as a high-level indicator of successful collaboration.
@@ 2.2 Individual Reflections and Collaboration
A central question in our study is how these individual perceptions of collaboration relate to the behavioral traces we can sense and log. We do not assume that users’ subjective responses can be directly averaged into a “true” group state. Instead, we examine the distribution of perceptions across the group: Are they aligned or divergent? Are certain individuals systematically disengaged or overloaded? These variations matter because they signal different interpretations of what collaboration felt like to the user. By examining the relationship between sensor-based indicators (§2.1.2) and subjective reports of internal individual perception (§2.1.1), we can start to ask: Do shared gaze events map onto users’ awareness of shared attention? Does higher proximity align with perceptions of a sense of group cohesion? Does a more balanced distribution of interactions align with users’ perception that they and their peers contributed meaningfully to the task?
Importantly, the subjective states are not outcomes of the sensor data; they are independent yet complementary accounts of collaboration. Our framework treats them as the interpretive anchor against which we examine the dynamics captured by the system. This allows us to understand what groups did and how collaboration was experienced in the future, yielding insight into how systems might better reflect, support, or adapt to group behavior in MR environments. Together, these indicators form a continuous stream of behavioral data interpreted as isolated features but as observable patterns of interaction that may or may not align with how users interpret collaboration subjectively. Our goal is not to classify groups as ”effective” or “cohesive” based on signals but to assess where and how the system’s behavioral observations reflect users’ own perspectives.

### v2

@@ Abstract
Mixed Reality (MR) enables rich, embodied collaboration, yet it is uncertain if sensor and system-logged behavioral signals capture how users experience that collaboration. This disconnect stems from a fundamental gap: behavioral signals are observable and continuous, while collaboration is interpreted subjectively, and shaped by internal states like presence, cognitive availability, and social awareness. Our core insight is that sensor signals serve as observable manifestations of subjective experiences in MR collaboration, and they can be captured through sensor data such as shared gaze, speech, spatial movement, and other system-logged performance metrics. We propose the Sensor-to-Subjective (S2S) Mapping Framework, a conceptual model that links observable interaction patterns to users’ subjective perceptions of collaboration and internal cognitive states through sensor-based indicators and task performance metrics. To validate this model, we conducted a study with 48 participants across 12 MR groups engaged in a collaborative image-sorting task. Our findings show a correlation between sensed behavior and perceived collaboration, particularly through shared attention and proximity.
@@
Mixed Reality, Collaboration, PerceptionIntroduction Collaboration in mixed reality (MR)111MR aligns with Milgram’s reality-virtuality continuum [38], closely related to Augmented Reality (AR), where virtual elements seamlessly integrate with the real world. environments are becoming increasingly prevalent across domains such as education, design, healthcare, and remote work [18]. MR uniquely blends physical and virtual realms, allowing co-located and remote users to share digital content seamlessly. As MR becomes more immersive and integral to collaboration, researchers have investigated the effects of gaze, gesture sharing, and virtual replicas on the presence and cognitive load [4, 62]. Despite these advances, a fundamental challenge persists: How can we effectively understand the quality of collaboration in MR?
Modern MR systems generate extensive behavioral data through system logs and embedded sensors, capturing where users look [47], how they move [34], and when they speak [52], yet these signals often provide only limited insights into users’ internal states [73]. Prior work shows that speech, gaze, and turn-taking are tightly connected to the internal perception of collaboration [11]. However, two core limitations remain: (1) it is unclear whether sensor data meaningfully reflects how participants perceive collaboration, or if key experiential dimensions go undetected. (2) Existing studies demonstrate specific bi-modal correlation mappings such as gaze to presence [4], and physiological synchrony to co-presence [9], but stop short of offering a holistic objective-to-subjective mapping of collaboration.
We address these gaps through a systematic, multimodal Sensor-to-Subjective (S2S) mapping framework, applied in a multi-perspective study of collaborative MR involving 48 participants across 12 teams. First, we analyze sensor-derived indicators of group interaction, including gaze-based shared attention, conversation dynamics through audio, and spatial proximity, which can be captured through embedded sensing in MR headsets [73, 70]. Second, we examine task performance using system logs, measuring timing, interaction patterns, and decision changes throughout the collaboration. Finally, we assess subjective experience through post-exposure questionnaires that capture participants’ perceptions of group dynamics as well as their individual sense of presence and cognitive effort.
@@ 1.2 Sensor-Based Indicators of Group Behavior
Sensor-based methods across ubiquitous and immersive computing characterize group behavior through multiple data streams (gaze, voice, motion, location) without relying on external observation or self-reporting. Remote collaboration systems, such as TeamSense [73] and CoCo [50] monitor nonverbal cues to track group cohesion, while wearable systems assess synchrony and proximity patterns during collaborative tasks [58]. In MR, headset-embedded sensors can track gaze, object manipulation, and gesture alignment, which researchers have shown correspond to collaboration quality [70, 26]. Studies have correlated gaze and gesture sharing with co-presence [4], and linked physiological synchrony and gaze with perceived collaboration quality [9]. Other systems enhance spatial awareness by embedding gaze cues [28], while multimodal integrations of gaze and gesture have been used to improve performance and presence [5]. Additional work shows high-accuracy cognitive load prediction using multimodal sensing [25] and demonstrates that even minimal AR cues can influence team experience and [61]. Conversation-focused research further emphasizes how speech alignment, gesture, gaze, and turn-taking synchronize to create effective collaboration in co-located settings [60, 11]. These findings emphasize that collaboration is fundamentally multimodal, involving a convergence of signals across time.
Despite these insights, existing work remains limited in two ways: (1) isolated pairwise mappings rather than producing a generalizable mapping model and (2) lens-like views often aim to enhance either performance outcomes or structural interaction properties (such as coordination patterns and role distribution) without examining their relationship to participants’ subjective experiences. Prior work has explicitly recognized the need for structured frameworks that integrate multiple sensing modalities and interpret internal collaboration states. Barrett et al. call for multi-method frameworks that integrate sensor modalities to capture internal collaboration states [48], while social signal processing emphasizes the need to fuse multimodal behavioral signals to capture fine-grained group dynamics [61]. Systematic reviews also note the absence of frameworks that link objective data to users’ subjective experiences [17]. Despite these advances, the critical question of whether sensor-based behavioral indicators meaningfully reflect users’ internal perceptions of collaboration remains largely unexplored, particularly in co-located MR tasks using commercial headsets.
@@ 1.3 Subjective Indicators of Group Behavior
Subjective self-report surveys remain central to understanding how people experience collaboration. Standardized tools, such as post-task surveys, presence questionnaires [68, 51], and cognitive load surveys [23], are widely used in the Human-Computer Interaction (HCI) to gauge how users evaluate, interpret, and internalize virtual experiences. Prior work has used these subjective measures to assess collaborative quality in various contexts, including team cohesion in distributed work [73], perceived social presence in remote collaboration [45], and mutual understanding during joint problem-solving [72]. Educational studies use them to evaluate contribution equity and group satisfaction [67], while collaborative design work leverages them to assess idea sharing and creative synergy [41].
Despite rich insights, these subjective measures are rarely systematically linked to sensor-based behavioral data. This disconnect raises key questions: can gaze, speech, and proximity signals in MR environments meaningfully reflect users’ internal sense of collaboration? Do behavioral patterns align with perception, or challenge our assumptions? Answering these questions requires bridging sensor data and subjective experience in collaborative MR.
@@ 2.1.1 Subjective Perception of Collaboration
We define collaboration as not just joint actions, but also how individuals interpret, feel, and reflect on their shared experience. In MR, where social cues, task progress, and attention are distributed across a blend of physical and virtual elements, shaped by both individual states and group-level awareness. We treat this as the subjective perception of collaboration and operationalize it through individual-level and group-level dimensions.
Individual-Level Experience. Each user’s interpretation of collaboration is filtered through their sense of presence [54], cognitive availability [56], and personal contribution [14]. These passive background states directly influence how users perceive and make sense of social interaction [45]. At its most foundational, collaboration begins with how individuals experience being part of a group in each moment. In immersive settings, this includes:
Cognitive Load: The mental effort to manage task demands affects users’ ability to perceive coordination signals. A taxed user may miss coordination cues, while a cognitively available one may detect attention shifts or conversational breaks [15]. Thus, we hypothesize that, cognitive load influences not only task performance [30] but also perceived collaboration quality [59].
Contribution Awareness: Users’ sense of how meaningfully others contributed serves as a proxy for perceived participation balance [13]. This acts as a bridge between personal effort and social perception. It links internal judgment with observed peer input and becomes especially salient in unstructured, role-free MR settings.
@@
Group-Level Reflection. MR system tracks each user individually and users form post-task judgments about shared experience [57], such as whether they attended to the same objects, communicated effectively, and collaborated cohesively. We capture this through four constructs:
Shared Attention Awareness captures whether users felt others focused on the same virtual elements, indicating shared intent and coordination [39].
Conversational Support reflects whether dialogue helped users understand (clarify) and contribute (invite participation), signaling cohesion and mutual support [21].
Proximity Impact assesses how physical closeness, via MR, affected collaboration responsiveness. Prior work suggests that co-located interaction, being physically near others, improves responsiveness [20].
Group Collaboration measures users’ overall sense of team effectiveness based on observed behaviors, internal judgments, and implicit comparison to collaborative expectations [19].
These individually collected measures reflect each user’s interpretation of the group, shaped by their state, engagement, and task involvement, forming subjective readings of an inherently social phenomenon.
@@ 2.1.2 Objective Perception of Collaboration
While subjective reports capture how collaboration feels, immersive systems log how it unfolds. Although these are individual-level measurements, they are not isolated; they reflect what each user does in relation to others in the group. The same user performing the task alone would not produce the same sensor traces. These sensor-based signals (derived from sensor data and system instrumentation) are shaped by interaction: who a user looks at, how close they are to others, and when they speak or remain silent. Interpreting such distributed traces in group contexts is key to understanding collaboration objectively. We focus on three core indicators supported by prior work [70, 73] and accessible on commodity MR headsets [37, 1, 2]:
Conversation and speech activity have long been used to study group behavior across psychology, anthropology, and HCI [33, 12]. Recent work using ubiquitous sensors underscores the continued importance of conversation [32]. In MR, where interaction blends physical and virtual contexts, speech activity offers valuable insight into collaboration [70]. To this end, logged via headset microphones, we extract speaking time, turn count, and participation variance (non-transcribing content). These metrics, known to reflect dominance and fluency [42], offer insight into how evenly members contribute.
Shared attention, the ability to jointly focus on an object, underpins coordination and social connection [69]. It has been widely studied in collaborative systems [8, 39], often through gaze-based detection[72]. We build on this by using MR headset eye-tracking to detect overlapping gaze on the same virtual object, enabling continuous, time-resolved logging of shared attention [39]. Though not explicitly communicated, such alignment reflects group awareness and may shape perceived collaboration. We aim to assess whether sensed shared attention reflects users’ perception of collaboration.
@@
Proximity influences collaboration by shaping knowledge exchange, creativity, and group cohesion [66]. In co-located collaborative scenarios, groups tend to perform better when members are physically near one another rather than dispersed [24, 65]. Studies have observed that individuals often move closer to those with whom they share stronger social bonds [16]. We capture proximity as the average pairwise distance between users using headset position data in a shared coordinate frame.
@@
We focus on group-level distributions of proximity and conversation (overall closeness or speech participation) rather than directional relationships (who was closest to whom or who spoke to whom). Our goal is to capture collective engagement, not dominant roles. Prior work shows that proximity itself (regardless of direction) signals social engagement [16], and evenly distributed proximity and turn-taking indicate group cohesion [73]. Moreover, our subjective measures focus on perceived group collaboration, which is more likely influenced by general closeness than specific spatial arrangements.
@@
Beyond sensor signals, we log task-relevant interactions such as how often each user interacts with content, how frequently prior actions are revised, and the number of unique configurations a group explores before reaching a consensus. These metrics reflect participation balance, coordination, and decision negotiation. We also record task completion time as a coarse measure of collaborative success.
@@ 2.2 Individual Reflections and Collaboration
A central question in our study is how individual perceptions of collaboration relate to the sensed behavioral traces. Rather than averaging subjective reports into a “single group state”, we examine the distribution of perceptions across the group: Are they aligned or divergent? Do some users feel disengaged or overloaded? Such differences reflect varied interpretations of what collaboration felt like to the user. By linking sensor-based indicators (§2.1.2) with subjective internal perceptions (§2.1.1), we ask: Do shared gaze events map onto users’ awareness of shared attention? Does proximity align with perceived cohesion? Do balanced interactions align with users’ perception that they and their peers contributed meaningfully to the task?
Importantly, the subjective states are not outcomes of the sensor data; they are independent yet complementary accounts of collaboration. Our framework treats them as the interpretive anchor against which we examine the dynamics captured by the system. This allows us to understand what groups did and how collaboration was experienced in the future, yielding insight into how systems might better reflect, support, or adapt to group behavior in MR environments. Together, these indicators form a continuous stream of behavioral data interpreted as isolated features but as observable patterns of interaction that may or may not align with how users interpret collaboration subjectively. Rather than labeling groups as “effective” or “cohesive,” our goal is to assess whether and how the system’s observations reflect how collaboration was actually experienced.